# AP2O-Coder: Human-Inspired Progressive Optimization to Fix LLM Code Errors

### v3

@@ Abstract
LLMs’ code generation capabilities have yielded substantial improvements in the effectiveness of programming tasks. However, LLM-generated code still suffers from compilation and runtime errors. Existing preference optimization methods primarily focus on enhancing LLMs’ coding abilities using pass/fail signals in the preference data, overlooking the deep-level error types in the failed codes. To address this, we propose Adaptively Progressive Preference Optimization (AP2O) for coding (i.e., AP2O-Coder), a method that guides LLMs adaptively and methodically to reduce code errors for code generation. Specifically, we construct an error notebook from failed codes and progressively optimize the LLM to correct errors type by type. Furthermore, we adaptively replay error types to tailor to the LLM’s evolving weaknesses throughout training. Through extensive experiments on both code and general LLMs (Llama, Qwen, and DeepSeek series) with parameters ranging from 0.5B to 34B, our AP2O-Coder improves code generation performance by up to 3% in $pass@k$ while using less preference data. Code: https://github.com/TsingZ0/AP2O.
@@ Introduction
Among all the capabilities of large language models (LLMs), code generation is one of the most attractive abilities (sheokand2025codemixbench; dou2024stepcoder). However, LLM-generated code still suffers from compilation and runtime errors (tambon2025bugs), such as SyntaxError and TypeError. Reinforcement Learning with Verifiable Rewards (RLVR) is a powerful technique for post-training to correct pre-trained LLMs’ weaknesses, particularly in the code domain (yue2025does; zhao2025absolute; wang2025reinforcement). It only requires the problem prompts and unit tests to construct training data, with no need for output answers (codes). The LLM can self-generate multiple answers for each problem and use the corresponding unit tests to verify the correctness of these answers, automatically obtaining pass/fail signals (liurltf).
@@
Nevertheless, online RL approaches are unstable during training due to the changing models or environments (moskovitz2023confrontingrewardmodeloveroptimization). As an offline method, Direct Preference Optimization (DPO) (rafailov2023direct) was introduced as a more stable alternative that does not require reward models and can be easily applied with verifiable rewards. However, DPO and its variants (liu2025survey; pattnaik2024enhancing; meng2024simpo; croitoru2025curriculum) with identical data utilization exhibit three key shortcomings in reducing self-generated code errors: (1) unawareness of code errors, as preference data is constructed solely from pass/fail signals; (2) inability to focus on specific error types, since errors appear randomly in each training batch; and (3) neglect of the LLM’s changing weaknesses, as DPO samples preference data only once, and the static training set is pre-constructed, failing to adapt to the LLM’s updating ability during training process.
@@
To address these issues, we propose Adaptively Progressive Preference Optimization (AP2O), which consists of progressive preference optimization and adaptive error replay modules. We integrate AP2O into the code LLM training and sandbox evaluation pipeline, creating our AP2O-Coder. Inspired by human error correction practices (xu2023importance), we treat the acquisition of only pass/fail signals in existing DPO-based methods as taking exams, akin to grading exam papers. In DPO-based methods, LLMs are guided to reduce failed answers solely based on these pass/fail signals, making it difficult for the model to understand why, where, and how it fails. Therefore, our AP2O-Coder first analyzes the failed answers in the exam using a programming-language-specific analyzer (e.g., a Python interpreter), acting as an expert. After analyzing, we organize the errors into an error notebook, ordered by error frequency (ascending or descending). To mimic human correction practices and enhance code error correction effectiveness, we correct errors type by type based on this error notebook within the progressive preference optimization module. During the correction process, as the LLM is updated at each training step, the previously collected error notebook may no longer fit its current weaknesses. To mitigate this, we introduce the adaptive error replay module, which periodically evaluates the LLM on a small validation set, akin to taking small quizzes. This process identifies the error types in the LLM’s current failed answers and replays these error types, enabling the LLM to better focus on and correct them.
@@
Through the systematic process of exam, analysis, correction, and quiz, our AP2O-Coder outperforms five state-of-the-art baselines by up to 3% in $pass@k$ on EvalPlus (liu2023your) and LiveCodeBench v6 (jain2024livecodebench). This improvement is achieved across code and general LLMs, including CodeLlama (roziere2023code), DeepSeek-Coder (guo2024deepseek), Qwen2.5-Coder (hui2024qwen2), Llama3 (grattafiori2024llama), Qwen2.5 (team2024qwen2), and Qwen3 (yang2025qwen3), with parameter sizes ranging from 0.5B to 34B. We also find that progressing from low-to-high (L2H) error frequency is better for small models (e.g., 0.5B), while high-to-low (H2L) progression is more effective for large models (e.g., 34B). Our AP2O-Coder also requires a smaller amount of preference data, thanks to its organized and adaptive data utilization.
@@
LLM is becoming an essential tool and valuable companion for programming tasks, especially with the rise of code generation capabilities (cursor2025; anthropicclaude2025). There are three main post-training techniques for code generation tasks: (1) instruction tuning (ma2024llamoco; weyssow2023exploring), (2) model distillation (chen2023personalized; sun2024enhancing), and (3) reinforcement learning (RL) (mu2024clarifygpt; gehring2025rlef). Instruction tuning is a foundational approach for post-training tasks (lai2025survey), but it heavily depends on high-cost expert-written annotations, such as problem-code pairs for code generation (wu2025inversecoder). Although model distillation mitigates this by leveraging existing high-performance models, it suffers from issues like error propagation and data leakage (lei2023comprehensive). RL with human feedback (RLHF (kirkunderstanding)) is another approach, though RLHF can be biased and conflicting (xiao2024algorithmic; cheng2024reinforcement). In contrast, RL with verifiable rewards (RLVR (zhao2025absolute)) has garnered increasing attention in recent years. While online RL suffers from instability caused by model and environment shifts during training (moskovitz2023confrontingrewardmodeloveroptimization), offline methods—especially offline preference optimization—offer greater stability when managing self-generated code errors. Hence, this paper focuses on offline preference optimization.
@@
There are a few offline preference optimization methods specifically proposed or evaluated for coding tasks (liu2025survey; da2025agent), with most research focused on mathematical tasks (liu2025trust). In this work, we review general offline preference optimization methods, particularly those related to data utilization. These can be categorized into two main approaches: (1) dynamic sampling (rao2025dynamic; gee2025code) and (2) curriculum learning (pattnaik2024enhancing; shi2025cudip; hou2025novo; li20252d). Dynamic sampling methods mainly focus on resampling or active learning. Most resampling approaches require complex data quality criteria (hu2024most) or auxiliary models (huang2025adaptive), limiting their utility. Active learning approaches (muldrew2024active; xia2025selection) typically require re-training after each data selection step or rely on external large oracle LLMs (e.g., GPT-4 (achiam2023gpt)), resulting in high costs. On the other hand, curriculum learning requires easy/hard task criteria (lin2025survey), which are often absent and difficult to establish, particularly in complex post-training tasks like correcting self-generated code errors.
@@
Among offline preference optimization methods, DPO (rafailov2023direct) is both fundamental and widely used, so we begin with it. DPO does not require a critic model and a reward model. Instead, it directly leverages the contrastive relationships among preference pairs from $\mathcal{P}$. Based on the Bradley-Terry model (bradley1952rank), preference probability model that $y_{w}$ is preferred over $y_{l}$ is
@@
To address these shortcomings, inspired by human error correction practices (yang2021testing), we propose AP2O-Coder for AI coding tasks to: (1) construct an error notebook by collecting and analyzing errors, (2) guide the LLM to focus on correcting errors type by type, and (3) adaptively adjust the focus through small quizzes to fit the LLM’s current capacity.
@@
Initially, to assess a given LLM’s baseline ability for subsequent targeted and personalized correction, we have the LLM take exams on $M$ coding problems $\{x^{m}\}_{m=1}^{M}$ and evaluate its answers using the corresponding multiple unit tests $\{(ut^{1},ut^{2},\ldots)^{m}\}_{m=1}^{M}$. Since it is difficult to gather a sufficient number of high-quality coding problems with unit tests, our AP2O-Coder allow the LLM $\theta$ to generate $N$ answers (using a high temperature value [^1] Following the widely used temperature setting for exploration (shao2024deepseekmath), our AP2O-Coder set it to 1.0.) for each problem to thoroughly explore the LLM’s capability limits. Subsequently, our AP2O-Coder obtains the grading results (pass or fail) for each problem, which serve as the intermediate LLM-generated preference data, denoted as $\mathcal{D}_{\rm tr}=\{(x_{,}y^{1}_{p},\ldots,y^{N_{p}}_{p},y^{1}_{f},\ldots,y^{N_{f}}_{f})^{m}\}^{M}_{m=1}$, where $p$ and $f$ are short for passed and failed, respectively, and $N_{p}+N_{f}=N$. Formally, we have
@@
Simply knowing whether an answer is correct or incorrect does not provide enough information for the LLM to improve itself, especially on complex tasks like code error correction. Inspired once again by human error correction practices (xu2023importance), we propose diagnosing failed answers through detailed error type analysis and organizing them into an error notebook. However, the challenge lies in the need for an analyzer (expert) to perform the error diagnosis.
@@
We evaluate the effectiveness of AP2O-Coder by applying it to popular, state-of-the-art (SOTA) open-sourced code LLMs and general LLMs (Instruct versions) and post-training them to improve code generation performance. Code LLMs: CodeLlama (roziere2023code), DeepSeek-Coder (guo2024deepseek), and Qwen2.5-Coder (hui2024qwen2). General LLMs: Llama3 (grattafiori2024llama), Qwen2.5 (team2024qwen2), and Qwen3 (yang2025qwen3). We use LLMs ranging from 0.5B to 34B parameters.
@@
Since AP2O-Coder operates as an offline preference optimization method that emphasizes progression through code preference data pairs, we select the following related baselines for comparison in the code domain. (1) Init: The initial pre-trained (code) LLMs; (2) SFT-Coder: Optimizing the pre-trained LLMs via supervised fine-tuning (dodge2020fine) on coding tasks; (3) DPO-Coder: Using DPO (rafailov2023direct) with code-domain-specific training and sandbox evaluation pipelines; (4) Curri-DPO-Coder (pattnaik2024enhancing): A representative curriculum DPO variant with code-specific pipelines; (5) Dyn-DPO-Coder (gee2025code): A DPO variant that replaces the static preference dataset with dynamically sampled preference data during training progress. As for our AP2O-Coder, we have two versions: AP2O-Coder (L2H), and AP2O-Coder (H2L), corresponding to two progression directions of the progressive preference optimization module.
@@
Here, we focus on Python, one of the most frequently used programming languages. To obtain LLM-generated preference data, we use the coding problems and unit tests from the training/validation sets of MBPP (austin2021program) (384/90 problems for training/validation) and TACO (li2023taco) (1678/420 problems for training/validation), respectively. We use MBPP by default. Since we focus on fine-grained learning from failed answers, we filter out coding problems with fewer than two failed answers. As the code answers are self-generated, the filter results are specific to the ability of the given LLMs but remains consistent across all baselines.
@@
Building on existing code LLM works (team2024qwen2; hui2024qwen2), we use popular benchmarks such as EvalPlus (liu2023your) and LiveCodeBench v6 (Feb 2025–Apr 2025) (jain2024livecodebench), evaluating them with two metrics: $pass@k$ ($k\in\{1,5,10\}$) (roziere2023code) and sample efficiency (gao2022sample) with a temperature 0.6. Here, sample efficiency refers to the amount of data required during post-training. We conduct three training trials and report the average values. For more details and results, please refer to the Appendix.
@@
In contrast to AP2O-Coder’s consistent improvements, all baselines occasionally degrade performance compared to the initial pre-trained models, particularly on large models. This is mainly due to the catastrophic forgetting problem (li2025analyzing), common in post-training methods (e.g., SFT and most DPO variants) (fernando2024mitigating), where continual optimization on a small post-training set leads to overfitting and loss of previously learned generalizable knowledge. Directly organizing the training data order, as done in Curri-DPO-Coder, also fails to mitigate this issue and may even exacerbate it, as shown in Table 2. Since AP2O-Coder assesses the current capabilities of the gradually updating LLM and adaptively replays failed answers that fit its present weaknesses during the quiz phase, we can recover previously learned generalizable knowledge while effectively acquiring new knowledge.
@@
Figure 2: The statistics of errors on the test benchmark using Qwen2.5-Coder-7B. Best viewed when zoomed in.
@@
We evaluate the generalization ability of post-trained code LLMs by benchmarking $pass@k$ for larger values of $k$ (i.e., $k\in{5,10}$), as shown in Figure 4. A commonly observed phenomenon in the literature is that post-training often improves $pass@1$ while degrading performance at higher $k$ values (yue2025does; lyu2025top). This is also evident in Figure 4, where Curri-DPO-Coder shows significant performance degradation on large models, indicating that it may exacerbate the catastrophic forgetting problem. In contrast, our AP2O-Coder (H2L) not only maintains improvements at $pass@1$ but also enhances generalization at larger $k$ values. This is attributed to its ability to emphasize low-frequency errors in the later stages of training.
@@
In addition to improvements in pass rate, our AP2O-Coder also demonstrates greater data efficiency, requiring only 4%–60% preference data pairs compared to the DPO’s requirements, which is especially evident on large models. The H2L variant is more data-efficient than L2H, as prioritizing correcting high-frequency errors aligns with efficient human learning strategies (larionova2022frequency). As shown in Figure 5, Curri-DPO-Coder exhibits the opposite trend, needing more data for larger models. Although Dyn-DPO-Coder uses the least data, its performance is poor, as shown in Table 1 and Table 2.
@@
The evaluations above focus on existing code LLMs. Here, we demonstrate that our AP2O-Coder can also effectively adapt pre-trained general LLMs to the code domain. Notably, some general models (e.g., Qwen3) are “thinking” models that tend to generate lengthy reasoning by default, leading to lower pass rates due to the 512-token budget constraint on the MBPP benchmark set by EvalPlus (liu2023your). While SFT struggles to mitigate this issue, the other offline preference optimization methods perform better in Figure 6. Among these, our AP2O-Coder (L2H) achieves the best results, as the general 7B+ LLMs are poor in the specific code domain and benefit more from low-to-high-frequency (L2H) progressive optimization.
@@
Following prior work (xiao2024cal), we illustrate the training dynamics of our AP2O-Coder in Figure 8. The results show that optimization converges, with the H2L version outperforming L2H in reward measurement.

### v4

# AP2O-Coder: Adaptively Progressive Preference Optimization for Reducing Compilation and Runtime Errors in LLM-Generated Code

@@ Abstract
LLMs’ code generation capabilities have yielded substantial improvements in the effectiveness of programming tasks. However, LLM-generated code still suffers from compilation and runtime errors. Existing preference optimization methods primarily focus on enhancing LLMs’ coding abilities using pass/fail signals in the preference data, overlooking the deep-level error types in the failed codes. To address this, we propose Adaptively Progressive Preference Optimization (AP2O) for coding (i.e., AP2O-Coder), a method that guides LLMs adaptively and methodically to reduce code errors for code generation. Specifically, we construct an error notebook from failed codes and progressively optimize the LLM to correct errors type by type. Furthermore, we adaptively replay error types to tailor to the LLM’s evolving weaknesses throughout training. Through extensive experiments on both code and general LLMs (Llama, Qwen, and DeepSeek series) with parameters ranging from 0.5B to 34B, our AP2O-Coder improves code generation performance by up to 3% in $pass@k$ while using less preference data.
Codehttps://github.com/TsingZ0/AP2O
@@ Introduction
Among all the capabilities of large language models (LLMs), code generation is one of the most attractive abilities (Sheokand and Sawant 2025; Dou et al. 2024). However, LLM-generated code still suffers from compilation and runtime errors (Tambon et al. 2025), such as SyntaxError and TypeError. Reinforcement Learning with Verifiable Rewards (RLVR) is a powerful technique for post-training to correct pre-trained LLMs’ weaknesses, particularly in the code domain (Yue et al. 2025; Zhao et al. 2025; Wang et al. 2025). It only requires the problem prompts and unit tests to construct training data, with no need for output answers (codes). The LLM can self-generate multiple answers for each problem and use the corresponding unit tests to verify the correctness of these answers, automatically obtaining pass/fail signals (Liu et al. 2024).
@@
Nevertheless, online RL approaches are unstable during training due to the changing models or environments (Moskovitz et al. 2024). As an offline method, Direct Preference Optimization (DPO) (Rafailov et al. 2023) was introduced as a more stable alternative that does not require reward models and can be easily applied with verifiable rewards. However, DPO and its variants (Liu et al. 2025a; Pattnaik et al. 2024; Meng et al. 2024; Croitoru et al. 2025) with identical data utilization exhibit three key shortcomings in reducing self-generated code errors: (1) unawareness of code errors, as preference data is constructed solely from pass/fail signals; (2) inability to focus on specific error types, since errors appear randomly in each training batch; and (3) neglect of the LLM’s changing weaknesses, as DPO samples preference data only once, and the static training set is pre-constructed, failing to adapt to the LLM’s updating ability during training process.
@@
To address these issues, we propose Adaptively Progressive Preference Optimization (AP2O), which consists of progressive preference optimization and adaptive error replay modules. We integrate AP2O into the code LLM training and sandbox evaluation pipeline, creating our AP2O-Coder. Inspired by human error correction practices (Xu 2023), we treat the acquisition of only pass/fail signals in existing DPO-based methods as taking exams, akin to grading exam papers. In DPO-based methods, LLMs are guided to reduce failed answers solely based on these pass/fail signals, making it difficult for the model to understand why, where, and how it fails. Therefore, our AP2O-Coder first analyzes the failed answers in the exam using a programming-language-specific analyzer (e.g., a Python interpreter), acting as an expert. After analyzing, we organize the errors into an error notebook, ordered by error frequency (ascending or descending). To mimic human correction practices and enhance code error correction effectiveness, we correct errors type by type based on this error notebook within the progressive preference optimization module. During the correction process, as the LLM is updated at each training step, the previously collected error notebook may no longer fit its current weaknesses. To mitigate this, we introduce the adaptive error replay module, which periodically evaluates the LLM on a small validation set, akin to taking small quizzes. This process identifies the error types in the LLM’s current failed answers and replays these error types, enabling the LLM to better focus on and correct them.
@@
Through the systematic process of exam, analysis, correction, and quiz, our AP2O-Coder outperforms five state-of-the-art baselines by up to 3% in $pass@k$ on EvalPlus (Liu et al. 2023) and LiveCodeBench v6 (Jain et al. 2024). This improvement is achieved across code and general LLMs, including CodeLlama (Roziere et al. 2023), DeepSeek-Coder (Guo et al. 2024), Qwen2.5-Coder (Hui et al. 2024), Llama3 (Grattafiori et al. 2024), Qwen2.5 (Qwen Team 2024), and Qwen3 (Yang et al. 2025), with parameter sizes ranging from 0.5B to 34B. We also find that progressing from low-to-high (L2H) error frequency is better for small models (e.g., 0.5B), while high-to-low (H2L) progression is more effective for large models (e.g., 34B). Our AP2O-Coder also requires a smaller amount of preference data, thanks to its organized and adaptive data utilization.
@@
LLM is becoming an essential tool and valuable companion for programming tasks, especially with the rise of code generation capabilities (Cursor 2025; Anthropic 2025). There are three main post-training techniques for code generation tasks: (1) instruction tuning (Ma et al. 2024; Weyssow et al. 2023), (2) model distillation (Chen et al. 2023; Sun et al. 2024), and (3) reinforcement learning (RL) (Mu et al. 2024; Gehring et al. 2025). Instruction tuning is a foundational approach for post-training tasks (Lai et al. 2025), but it heavily depends on high-cost expert-written annotations, such as problem-code pairs for code generation (Wu et al. 2025). Although model distillation mitigates this by leveraging existing high-performance models, it suffers from issues like error propagation and data leakage (Lei and Tao 2023). RL with human feedback (RLHF (Kirk et al. 2024)) is another approach, though RLHF can be biased and conflicting (Xiao et al. 2024a; Cheng et al. 2024). In contrast, RL with verifiable rewards (RLVR (Zhao et al. 2025)) has garnered increasing attention in recent years. While online RL suffers from instability caused by model and environment shifts during training (Moskovitz et al. 2024), offline methods—especially offline preference optimization—offer greater stability when managing self-generated code errors. Hence, this paper focuses on offline preference optimization.
@@
There are a few offline preference optimization methods specifically proposed or evaluated for coding tasks (Liu et al. 2025a; Da et al. 2025), with most research focused on mathematical tasks (Liu et al. 2025b). In this work, we review general offline preference optimization methods, particularly those related to data utilization. These can be categorized into two main approaches: (1) dynamic sampling (Rao et al. 2025; Gee et al. 2025) and (2) curriculum learning (Pattnaik et al. 2024; Shi et al. 2025; Hou 2025; Li and Zhang 2025). Dynamic sampling methods mainly focus on resampling or active learning. Most resampling approaches require complex data quality criteria (Hu et al. 2024) or auxiliary models (Huang et al. 2025), limiting their utility. Active learning approaches (Muldrew et al. 2024; Xia et al. 2025) typically require re-training after each data selection step or rely on external large oracle LLMs (e.g., GPT-4 (Achiam et al. 2023)), resulting in high costs. On the other hand, curriculum learning requires easy/hard task criteria (Lin et al. 2025), which are often absent and difficult to establish, particularly in complex post-training tasks like correcting self-generated code errors.
@@
Among offline preference optimization methods, DPO (Rafailov et al. 2023) is both fundamental and widely used, so we begin with it. DPO does not require a critic model and a reward model. Instead, it directly leverages the contrastive relationships among preference pairs from $\mathcal{P}$. Based on the Bradley-Terry model (Bradley and Terry 1952), preference probability model that $y_{w}$ is preferred over $y_{l}$ is
@@
To address these shortcomings, inspired by human error correction practices (Yang et al. 2021), we propose AP2O-Coder for AI coding tasks to: (1) construct an error notebook by collecting and analyzing errors, (2) guide the LLM to focus on correcting errors type by type, and (3) adaptively adjust the focus through small quizzes to fit the LLM’s current capacity.
@@
Initially, to assess a given LLM’s baseline ability for subsequent targeted and personalized correction, we have the LLM take exams on $M$ coding problems $\{x^{m}\}_{m=1}^{M}$ and evaluate its answers using the corresponding multiple unit tests $\{(ut^{1},ut^{2},\ldots)^{m}\}_{m=1}^{M}$. Since it is difficult to gather a sufficient number of high-quality coding problems with unit tests, our AP2O-Coder allow the LLM $\theta$ to generate $N$ answers (using a high temperature value [^1] Following the widely used temperature setting for exploration (Shao et al. 2024), our AP2O-Coder set it to 1.0.) for each problem to thoroughly explore the LLM’s capability limits. Subsequently, our AP2O-Coder obtains the grading results (pass or fail) for each problem, which serve as the intermediate LLM-generated preference data, denoted as $\mathcal{D}_{\rm tr}=\{(x_{,}y^{1}_{p},\ldots,y^{N_{p}}_{p},y^{1}_{f},\ldots,y^{N_{f}}_{f})^{m}\}^{M}_{m=1}$, where $p$ and $f$ are short for passed and failed, respectively, and $N_{p}+N_{f}=N$. Formally, we have
@@
Simply knowing whether an answer is correct or incorrect does not provide enough information for the LLM to improve itself, especially on complex tasks like code error correction. Inspired once again by human error correction practices (Xu 2023), we propose diagnosing failed answers through detailed error type analysis and organizing them into an error notebook. However, the challenge lies in the need for an analyzer (expert) to perform the error diagnosis.
@@
We evaluate the effectiveness of AP2O-Coder by applying it to popular, state-of-the-art (SOTA) open-sourced code LLMs and general LLMs (Instruct versions) and post-training them to improve code generation performance. Code LLMs: CodeLlama (Roziere et al. 2023), DeepSeek-Coder (Guo et al. 2024), and Qwen2.5-Coder (Hui et al. 2024). General LLMs: Llama3 (Grattafiori et al. 2024), Qwen2.5 (Qwen Team 2024), and Qwen3 (Yang et al. 2025). We use LLMs ranging from 0.5B to 34B parameters.
@@
Since AP2O-Coder operates as an offline preference optimization method that emphasizes progression through code preference data pairs, we select the following related baselines for comparison in the code domain. (1) Init: The initial pre-trained (code) LLMs; (2) SFT-Coder: Optimizing the pre-trained LLMs via supervised fine-tuning (Dodge et al. 2020) on coding tasks; (3) DPO-Coder: Using DPO (Rafailov et al. 2023) with code-domain-specific training and sandbox evaluation pipelines; (4) Curri-DPO-Coder (Pattnaik et al. 2024): A representative curriculum DPO variant with code-specific pipelines; (5) Dyn-DPO-Coder (Gee et al. 2025): A DPO variant that replaces the static preference dataset with dynamically sampled preference data during training progress. As for our AP2O-Coder, we have two versions: AP2O-Coder (L2H), and AP2O-Coder (H2L), corresponding to two progression directions of the progressive preference optimization module.
@@
Here, we focus on Python, one of the most frequently used programming languages. To obtain LLM-generated preference data, we use the coding problems and unit tests from the training/validation sets of MBPP (Austin et al. 2021) (384/90 problems for training/validation) and TACO (Li et al. 2023) (1678/420 problems for training/validation), respectively. We use MBPP by default. Since we focus on fine-grained learning from failed answers, we filter out coding problems with fewer than two failed answers. As the code answers are self-generated, the filter results are specific to the ability of the given LLMs but remains consistent across all baselines.
@@
Building on existing code LLM works (Qwen Team 2024; Hui et al. 2024), we use popular benchmarks such as EvalPlus (Liu et al. 2023) and LiveCodeBench v6 (Feb 2025–Apr 2025) (Jain et al. 2024), evaluating them with two metrics: $pass@k$ ($k\in\{1,5,10\}$) (Roziere et al. 2023) and sample efficiency (Gao et al. 2022) with a temperature 0.6. Here, sample efficiency refers to the amount of data required during post-training. We conduct three training trials and report the average values. For more details and results, please refer to the Appendix.
@@
In contrast to AP2O-Coder’s consistent improvements, all baselines occasionally degrade performance compared to the initial pre-trained models, particularly on large models. This is mainly due to the catastrophic forgetting problem (Li et al. 2025), common in post-training methods (e.g., SFT and most DPO variants) (Fernando et al. 2024), where continual optimization on a small post-training set leads to overfitting and loss of previously learned generalizable knowledge. Directly organizing the training data order, as done in Curri-DPO-Coder, also fails to mitigate this issue and may even exacerbate it, as shown in Table 2. Since AP2O-Coder assesses the current capabilities of the gradually updating LLM and adaptively replays failed answers that fit its present weaknesses during the quiz phase, we can recover previously learned generalizable knowledge while effectively acquiring new knowledge.
@@
Figure 2: The statistics of errors on the test benchmark using Qwen2.5-Coder-7B. “Er.” is short for “Error”.
@@
We evaluate the generalization ability of post-trained code LLMs by benchmarking $pass@k$ for larger values of $k$ (i.e., $k\in{5,10}$), as shown in Figure 4. A commonly observed phenomenon in the literature is that post-training often improves $pass@1$ while degrading performance at higher $k$ values (Yue et al. 2025; Lyu et al. 2025). This is also evident in Figure 4, where Curri-DPO-Coder shows significant performance degradation on large models, indicating that it may exacerbate the catastrophic forgetting problem. In contrast, our AP2O-Coder (H2L) not only maintains improvements at $pass@1$ but also enhances generalization at larger $k$ values. This is attributed to its ability to emphasize low-frequency errors in the later stages of training.
@@
In addition to improvements in pass rate, our AP2O-Coder also demonstrates greater data efficiency, requiring only 4%–60% preference data pairs compared to the DPO’s requirements, which is especially evident on large models. The H2L variant is more data-efficient than L2H, as prioritizing correcting high-frequency errors aligns with efficient human learning strategies (Larionova and Martynova 2022). As shown in Figure 5, Curri-DPO-Coder exhibits the opposite trend, needing more data for larger models. Although Dyn-DPO-Coder uses the least data, its performance is poor, as shown in Table 1 and Table 2.
@@
The evaluations above focus on existing code LLMs. Here, we demonstrate that our AP2O-Coder can also effectively adapt pre-trained general LLMs to the code domain. Notably, some general models (e.g., Qwen3) are “thinking” models that tend to generate lengthy reasoning by default, leading to lower pass rates due to the 512-token budget constraint on the MBPP benchmark set by EvalPlus (Liu et al. 2023). While SFT struggles to mitigate this issue, the other offline preference optimization methods perform better in Figure 6. Among these, our AP2O-Coder (L2H) achieves the best results, as the general 7B+ LLMs are poor in the specific code domain and benefit more from low-to-high-frequency (L2H) progressive optimization.
@@
Following prior work (Xiao et al. 2024b), we illustrate the training dynamics of our AP2O-Coder in Figure 8. The results show that optimization converges, with the H2L version outperforming L2H in reward measurement.